OpenStudy (anonymous):

i'm attempting to understand 8th grade math. Is there somewhere that i can learn how this equation can be answered? 2x2+3x=20

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

u factor it.

OpenStudy (anonymous):

u bring 20 to the other side of the equation to make it 2x^2+3x-20=0. then u need two numbers who's product is -20, but add up to 3.

OpenStudy (anonymous):

Yes! Multiply 2x2. You get 4. 4+3x = 20. Isolate the x. Subtract 4 on both sides. 3x =16 x = 16/3.

OpenStudy (anonymous):

so find factors of -20 that also add up to 3.

OpenStudy (anonymous):

still baffeled. this is muliple choice and the answer you give makes some sense but my options are 4, -2/5 or 4, -5/2 or -5, 5/2 or -4, 2/5. I want to understand but i'm an old guy trying to help my kid and its not making sense.

OpenStudy (anonymous):

ok, look. so the answer would look something like this: (2x-5)(x+4).

OpenStudy (anonymous):

all that equals to zero. (2x-5)(x+4)=0

OpenStudy (anonymous):

and then, you would use one quantity and make it equal to zero. (2x-5)=0 2x=5 x=5/2 This is one solution. x+4=0 x=-4. This is the second solution. Do you understand it at all? If not, let me know which part you do not understand.

OpenStudy (anonymous):

you have been very helpful but this is beyond my abilities. i cannot even comprehend the smallest amount of the equation. it's been 36 years since i have done anything even close to this. my daughter will need to go to school early or during lunch for help. thanks very much

OpenStudy (anonymous):

I can explain it if you would like me to. That is why I am here. :)

OpenStudy (anonymous):

i'll put my daughter at the keyboard

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

ok, so, I got (2x-5)(x+4)=0. If you use the distributive property, you can distribute each number and get: 2x^2+3x-20=0.

OpenStudy (anonymous):

okay so what is the first step and what are you solving for

OpenStudy (anonymous):

U r solving for x. That is, two solutions of x. Since you cannot add 2x^2 and 2x because they are not common terms, you would have to divide them to make them common terms. You know the simple operations of variables right? like x*x=x^2 and x+x=2x?

OpenStudy (anonymous):

yes i understand that but how do you know what the answer is ?

OpenStudy (anonymous):

you would have to put the variable x in two quantities, so that if you multiply using the distributive property, you would get the original equation, and that also, equal to 0. If you substitute two values (answers) into the original equation, it should give you an answer of 0, which is what the original equation is equal to.

OpenStudy (anonymous):

she went to a neighbor that helps his grandson. thanks again

OpenStudy (anonymous):

awww, she ditched me :( lol it's ok, it's kind of hard to teach online on chat anyway. I tried my best!